# Forward modeling fluctuations in the DESI LRGs target sample using image simulations

### v1

@@ Abstract
We use the forward modeling pipeline, Obiwan, to study the imaging systematics of the Luminous Red Galaxies (LRGs) targeted by the Dark Energy Spectroscopic Instrument (DESI). Imaging systematics refers to the false fluctuation of galaxy densities due to varying observing conditions and astrophysical foregrounds corresponding to the imaging surveys from which DESI LRG target galaxies are selected. We update the Obiwan pipeline, which had previously been developed to simulate the optical images used to target DESI data, to further simulate WISE images in the infrared. This addition makes it possible to simulate the DESI LRGs sample, which utilizes WISE data in the target selection. Deep DESI imaging data combined with a method to account for biases in their shapes is used to define a truth sample of potential LRG targets. We inject these data evenly throughout the DESI Legacy Imaging Survey footprint at declinations between -30 and 32.375 degrees. We simulate a total of 15 million galaxies to obtain a simulated LRG sample (Obiwan LRGs) that predicts the variations in target density due to imaging properties. We find that the simulations predict the trends with depth observed in the data, including how they depend on the intrinsic brightness of the galaxies. We observe that faint LRGs are the main contributing power of the imaging systematics trend induced by depth. We also find significant trends in the data against Galactic extinction that are not predicted by Obiwan. These trends depend strongly on the particular map of Galactic extinction chosen to test against, implying Large-Scale Structure systematic contamination (e.g. Cosmic-Infrared Background) in the Galactic extinction maps is a likely root cause. We additionally observe that the DESI LRGs sample exhibits a complex dependency on a combination of seeing, depth, and intrinsic galaxy brightness, which is not replicated by Obiwan, suggesting discrepancies between the current simulation settings and the actual observations. The detailed findings we present should be used to guide any observational systematics mitigation treatment for the clustering of the DESI LRG sample.
@@ 1 Introduction
Modern cosmology is built upon detecting signals from the distant universe. We collect light and gravity waves through various instruments at different frequencies [1] [2] [3] [4] [5]. For galaxy imaging surveys [6] [7] [8], in particular, we collect light profiles of galaxies by taking pictures of the night sky. These images help us infer the properties of galaxies, which can be used as fundamental input for model development in Cosmology. However, as light travels through the Milky Way and the Earth’s atmosphere to our camera, it is perturbed by different effects. These effects alter the light profile, making the galaxies in the images deviate from their true appearance. We call these effects imaging systematics [9] [10] [11] [12]. It is an umbrella term containing any possible effects that contribute to the false fluctuation of galaxy densities. In practice, we use survey property maps to quantify the strength of different effects on the galaxy densities. These maps include astronomical effects like Galactic dust extinction, stellar density (including stellar streams), as well as instrument-based effects like image background noise level and seeing. The correlation between these property maps and the galaxy density indicates how our galaxy samples are contaminated by each survey property.
@@
An accurate correction of imaging systematics lays a solid foundation for a stringent and unbiased measurement for all cosmological probes. It is particularly important for probes that are sensitive to the power spectrum. For example, primordial non-Gaussianity [13] [14] [15] is sensitive to the large-scale part of the power spectrum. Redshift Space distortion measurements [16] are sensitive to the overall shape. If not treated properly, imaging systematics would leave an imprint on the galaxy density map that would be degenerate with the primordial non-Gaussianity signal. On the other hand, traditional probes like Baryon Acoustic Oscillations [17] have a characteristic bump when computing their correlation function, thus are less sensitive to imaging systematics; it is possible to obtain a robust measurement even without any treatment of imaging systematics.
@@
As ongoing and future galaxy surveys yield increasingly precise data, the margin for error in imaging systematics needs to diminish to achieve a greater constraining power. The Dark Energy Spectroscopic Survey (DESI) [18] [19] [20] [21] [22] [23] [24] [25] [26] is a Stage-4 [27] cosmology survey that aims at obtaining better cosmological constraints than the former surveys [28] [29]. DESI is conducted on the Mayall 4-meter telescope at Kitt Peak National Observatory. It is designed to build an unprecedented 3D map of the universe. It will obtain 47 million galaxy and star spectra, and is expected to achieve a precision of smaller than 0.4% precision measurement of the BAO distance scale, and 1.05% precision on the Hubble parameter [30].
@@
Imaging systematics arises from varying observing conditions in imaging surveys. In the case of DESI, the galaxies are selected from the 9th data release (DR9) of the DESI Legacy Imaging Surveys [31] [32] [33]. The survey is comprised of 3 imaging projects on different telescopes: The Beijing-Arizona Sky Survey (BASS) [34], the DECam Legacy Survey (DECaLS) [35], and the Mayall z-band Legacy Survey (MzLS) [36]. These images are processed with the DESI image reduction pipeline Legacypipe [^1] https://github.com/legacysurvey/legacypipe. Legacypipe transforms these images into a catalog of astronomical objects.
@@
We developed a forward modeling pipeline, Obiwan [^2] https://github.com/DriftingPig/obiwan_code/tree/ObiwanLrgPaper, to study the effects of imaging systematics. The idea is to add simulated galaxies into real images, and extract these simulated galaxies with the image reduction pipeline Legacypipe [37] [38] [39]. This process traces the un-cleaned residuals in the image and provides a catalog that has the same imaging systematics variation as the tracers that we are interested in. Obiwan was first utilized to study the imaging systematics of eBOSS ELGs [40] [41], and it is further developed to fit our purpose here. Obiwan is modified from Legacypipe. The majority of the software modules remain untouched to ensure that we mimic this pixel-to-catalog transformation process.
Similar pipelines also exist in other surveys like the Dark Energy Survey (DES) [42] and the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) [^3] https://github.com/lsst/source_injection, and they follow the same philosophy. The biggest difference between these pipelines is which image reduction pipeline it is based on. Though this method has made a significant contribution to cosmological analysis, it still suffers from not being able to fully mimic the behavior of true galaxies, resulting in errors in scientific analysis like measuring photometric redshift, magnification, etc. We investigate the origin of the mismatch between real and simulated galaxies, and our conclusions can serve as a reference to make future improvements for studies based on similar samples.
@@
In this work, we use Obiwan to study the imaging systematics trend of DESI LRGs. We recover similar systematics trends on simulated LRGs that we call Obiwan LRGs. We also identify spurious trends that are possibly due to imperfect survey property maps or uncertainties in the image calibration stage. The results we present allow a deeper understanding of the sources of artificial variation in the density of DESI LRGs, which we expect will influence how this variation is corrected for in the ultimate clustering analyses of DESI LRGs.
@@
This paper is organized as follows. In section 2 we describe the data we used in our analysis. In section 3 we describe the pipeline, as well as new updates. In section 4 we describe our method to determine the target selection cut. In section 5 we discuss the process needed before simulation runs. In section 6 we discuss the analysis of output products. In section 7 we discuss the indication of our results.
@@ 2 Data
We use three datasets for our work. These datasets originate from the images from the DR9 DESI Legacy Imaging Survey and only the image selection criteria is different. The images are compiled from 4 cameras: The Dark Energy Camera (DECam, [43]) on the Victor M. Blanco 4-meter Telescope at the Cerro Tololo Inter-American Observatory (CTIO); the MOSAIC-3 camera [44], formerly mounted on the Nicholas U. Mayall 4-meter Telescope at the Kitt Peak National Observatory; the 90Prime camera [45] on the Bok 2.3-meter telescope located on Kitt Peak; and the Wide-field Infrared Survey Explorer (WISE, [46]) satellite.
The three datasets we described below are composed of these images and are processed with the pipeline Legacypipe. The DESI Legacy Imaging Survey DR9 described in section 2.1 uses most of images, excluding a few defective images. There are also images excluded because the selected region has reached the depth requirement. This set is used for mass production of Obiwan LRGs to study imaging systematics. The COSMOS Deep catalogs in section 2.3 utilize the regions of sky with a lot more exposures than a typical region, and has much deeper photometry. This set is used to select the truth sample that Obiwan injects into the images. The COSMOS Repeats in section 2.2 uses the large number of observations of the COSMOS field. The images are separated into 10 distinct sets and each set is individually processed. These repeats are used to understand photometric scattering on DESI LRGs to define extended color cut for the injected Obiwan LRG-like sample.
@@ 2.1 DESI Legacy Imaging Survey DR9
We use and simulate data from the DR9 DESI Legacy Imaging Survey [^4] https://www.legacysurvey.org/dr9/description/ [31], which was used to obtain the sample that DESI uses for target selection and follow-up spectroscopy [47] [48] [49] [50] [51] [52] [53] [54] [55]. We briefly describe these images here. This imaging survey consists of three optical bands: $g$, $r$, and $z$, and 4 infrared bands from WISE, $W1$, $W2$, $W3$, and $W4$. The optical band is separated into two parts: the South and the North, separated by declination (Dec) $32.375$ in the North Galactic Cap. The South part consists of exposures taken from DECam [56], with a typical seeing less than 1.3". The North part consists of images from the Beijing-Arizona Sky Survey [57] using the 90Prime camera for the $g$ and $r$ bands, and the Mosaic z-band Legacy Survey (MzLS) [58] for the $z$-band. The BASS survey has a typical seeing of less than 1.7" and MzLS has a seeing of less than 1.3". The pipeline transforms these images into a common grid with a pixel scale of 0.262 arcsec. These pixels are grouped into $3600\times 3600$ ‘bricks’ and imaging processing is performed at the individual brick level.
@@
The infrared bands are imaged by the WISE survey [59]. WISE is a space telescope, its orbits intersect with its Ecliptic poles ($|\textrm{Dec}|\approx 66$) while taking snapshots of the sky during its journey. Because of this orbiting plan, the WISE images have stable orientation in regions far away from the Ecliptic poles. For regions close to the Ecliptic poles, image orientation can be very different during different visits. Because of this difference, it is much easier to measure regions far from the Ecliptic poles: The point spread function (PSF) of WISE galaxies is a stable image, and its profile is determined by the orientation of the WISE telescope [60]. Our work focuses only on the Southern part of the imaging survey because we do not have truth input sample in the northern imaging area, and more details are explained in section 2.3. Though WISE is deeper near the Ecliptic poles, these regions are more difficult to model due to the different image orientations in different visits; however, we avoid these problematic regions since our footprint does not cover that region. The WISE survey has 4 bandpasses, but only 2 bands ($W1$ and $W2$) are used in DESI target selection, so we will focus on only $W1$ and $W2$ in what follows. These two bands have a typical point-spread function width of 6.5". The WISE images we process are the unWISE coadded data products [61, 62], which use a 2.75" pixel scale. Figure 1 is a representative source that has typical PSF size in all bands. It is located in a region that is observed by all 3 cameras in the optical bands. The first column shows what the real images look like in optical and infrared bands. The second column shows galaxy models of the real images. The third column shows the galaxy stamp for the simulated galaxy. The fourth column is the “blob map”: sources within this map are treated as blended sources, and will be fit simultaneously.
@@
Figure 1: Simulated targets in different bands and cameras, processed through the Legacypipe image reduction pipeline. The injected galaxy is located at the center of each image. The figures in the first column are the co-added images. The second column displays the galaxy/star profiles derived from fitting the image. The third column shows the simulated source that was injected into the images after convolution with the local PSF. The fourth column shows the contiguous area (known as a ‘blob’), where optical sources within the same blob are fit simultaneously (e.g., in order to resolve blending). The blobs are not used for WISE processing (see text for details) and thus not shown for their images. The top rows show the WISE $W1$ and $W2$ band images, respectively. The third row shows the DECam image in the $g$- band, and the forth row shows the $r$- band image from the BASS survey. The third row shows the $z$- band image from the MzLS survey.
@@ 2.2 COSMOS Repeats dataset
The COSMOS region has been imaged extensively by many instruments, including DECam. The COSMOS Repeats dataset we use is a set of catalogs covering $1\deg^{2}$ around (RA, Dec) = ($150.1\deg,2.2\deg$). This small region has a large number of exposures taken in all optical bands, which provides the ability to produce many independent subsets of data processed through Legacypipe and containing the same photometric sources. Ten such realizations were produced for the COSMOS repeats data. Gaussian noise was added to each image in each set so that the total galdepth [^5] It is defined as detection sensitivity of a round exponential galaxy with a radius of 0.45 arcsecond. in each set was equal to the DESI imaging requirements. The different image sets are grouped by seeing. The ten sets are labeled “set 0” through “set 9”, going from smallest to largest seeing (psfsize). The scatter in photometric measurements of sources common to each set of the COSMOS Repeats data thus provides a fully empirical determination of the expected variance in Legacypipe photometry in different seeing conditions. We simulate sources in the same sets of images with the same settings as COSMOS repeats catalog. We obtained 10 sets of corresponding simulated Obiwan COSMOS repeat catalogs that mimic the scattering behavior of sources in real COSMOS repeat catalogs. Achieving consistency with the distributions recovered from the real data in the COSMOS Repeats is our main method of validating the LRG-Obiwan pipeline. The COSMOS Repeats data will also be used to determine our target selection for the extended color boundary that we draw our truth sample from and will be discussed in Section 4. It will also be used to study PSF error contributions in section 7.3.

### v2

@@ Abstract
We use the forward modeling pipeline, Obiwan, to study the imaging systematics of the Luminous Red Galaxies (LRGs) targeted by the Dark Energy Spectroscopic Instrument (DESI). Imaging systematics refers to the false fluctuation of galaxy densities due to varying observing conditions and astrophysical foregrounds corresponding to the imaging surveys from which DESI LRG target galaxies are selected. We update the Obiwan pipeline, which we previously developed to simulate the optical images used to target DESI data, to further simulate WISE images in the infrared. This addition allows simulating the DESI LRGs sample, which utilizes WISE data in the target selection. Deep DESI imaging data combined with a method to account for biases in their shapes is used to define a truth sample of potential LRG targets. We inject these data evenly throughout the DESI Legacy Imaging Survey footprint at declinations between -30 and 32.375 degrees. We simulate a total of 15 million galaxies to obtain a simulated LRG sample (Obiwan LRGs) that predicts the variations in target density due to imaging properties. We find that the simulations predict the trends with depth observed in the data, including how they depend on the intrinsic brightness of the galaxies. We observe that faint LRGs are the main contributing source of the imaging systematics trend induced by depth. We also find significant trends in the data against Galactic extinction that are not predicted by Obiwan. These trends depend strongly on the particular map of Galactic extinction chosen to test against, implying systematic contamination in the Galactic extinction maps is a likely root cause (e.g., Cosmic-Infrared Background, dust temperature correction). We additionally observe a morphological change of the DESI LRGs population evidenced by a correlation between OII emission line average intensity and the size of the $z$-band PSF. This effect most likely results from uncertainties in background subtraction. The detailed findings we present should be used to guide any observational systematics mitigation treatment for the clustering of the DESI LRGs sample.
@@ 1 Introduction
Modern cosmology is built upon detecting signals from the distant universe. We collect light and gravity waves through various instruments at different frequencies [1, 2, 3, 4, 5]. For galaxy imaging surveys [6, 7, 8], in particular, we collect light profiles of galaxies by taking pictures of the night sky. These images help us infer the properties of galaxies, which can be used as fundamental input for model development in Cosmology. However, as light travels through the Milky Way and the Earth’s atmosphere to our camera, it is perturbed by different effects. These effects alter the light profile, making the galaxies in the images deviate from their true appearance. We call these effects imaging systematics [9, 10, 11, 12]. It is an umbrella term containing any possible effects contributing to galaxy densities’ false fluctuation. In practice, we use survey property maps to quantify the strength of different effects on the galaxy densities. These maps include astronomical effects like Galactic dust extinction and stellar density (including stellar streams) and instrument-based effects like image background noise level and seeing.
The correlation between these property maps and the galaxy density indicates how each survey property contaminates our galaxy samples.
@@
An accurate correction of imaging systematics lays a solid foundation for a stringent and unbiased measurement for all cosmological probes. It is essential for probes that are sensitive to the power spectrum. For example, primordial non-Gaussianity [13, 14, 15] is sensitive to the large-scale part of the power spectrum. Redshift Space distortion measurements [16] and full shape analysis [17, 18] are sensitive to the overall shape. If not treated properly, imaging systematics would leave an imprint on the galaxy density map that would be degenerate with the primordial non-Gaussianity signal. On the other hand, traditional probes like Baryon Acoustic Oscillations [19] have a characteristic bump when computing their correlation function, thus are less sensitive to imaging systematics; it is possible to obtain a robust measurement even without any treatment of imaging systematics.
@@
As ongoing and future galaxy surveys yield increasingly precise data, the margin for error in imaging systematics needs to diminish to achieve greater constraining power. The Dark Energy Spectroscopic Survey (DESI) [20, 21, 22, 23, 24, 25, 26, 27, 28] is a Stage-4 [29] cosmology survey that aims at obtaining better cosmological constraints than the former surveys [30, 31]. DESI is conducted on the Mayall 4-meter telescope at Kitt Peak National Observatory. It is designed to build an unprecedented 3D map of the universe. It will obtain 47 million galaxy and star spectra, and is expected to achieve a precision of less than 0.4% precision measurement of the BAO distance scale, and 1.05% precision on the Hubble parameter [32].
@@
Imaging systematics arises from varying observing conditions in imaging surveys. In the case of DESI, the galaxies are selected from the 9th data release (DR9) of the DESI Legacy Imaging Surveys [33, 34, 35]. The survey is comprised of 3 imaging projects on different telescopes: The Beijing-Arizona Sky Survey (BASS) [36], the DECam Legacy Survey (DECaLS) [37], and the Mayall z-band Legacy Survey (MzLS) [38]. These images are processed with the DESI image reduction pipeline Legacypipe [^1] https://github.com/legacysurvey/legacypipe. Legacypipe transforms these images into a catalog of astronomical objects.
@@
We developed a forward modeling pipeline, Obiwan [^2] https://github.com/DriftingPig/obiwan_code/tree/ObiwanLrgPaper, to study the effects of imaging systematics. The idea is to add simulated galaxies into real images and extract these simulated galaxies with the image reduction pipeline Legacypipe [39, 40, 41]. The pixel-to-object transformation is performed in the same way as real DESI galaxies. Obiwan traces the un-cleaned residuals in the image and provides a catalog with the same imaging systematics variation as the tracers we are interested in. Obiwan was first utilized to study the imaging systematics of eBOSS ELGs [42, 43], and it is further developed to fit our purpose here.
@@
In this work, we use Obiwan to study the imaging systematics trend of DESI LRGs[44]: The bright and luminous galaxies selected from DESI Legacy Imaging Surveys. They will be targeted to get spectrum throughout the 5-year DESI main survey. We recover similar systematics trends on simulated LRGs that we call Obiwan LRGs. We also identify spurious trends that are possibly due to imperfect survey property maps or uncertainties in the image calibration stage. The results we present allow a deeper understanding of the sources of artificial variation in the density of DESI LRGs, which we expect will influence how this variation is corrected for in the ultimate clustering analyses of DESI LRGs.
@@
This paper is organized as follows. In Section 2, we describe the data we used in our analysis. In Section 3 we describe the pipeline, as well as new updates. Section 4 describes our method to determine the target selection cut. In Section 5, we discuss the process needed before the simulation runs. In Section 6, we discuss the analysis of output products. In Section 7, we discuss the indication of our results.
@@ 2 Data
We use three datasets for our work. These datasets originate from the DR9 DESI Legacy Imaging Survey images, and only the image selection criteria are different. The images are compiled from 4 cameras: The Dark Energy Camera (DECam, [45]) on the Victor M. Blanco 4-meter Telescope at the Cerro Tololo Inter-American Observatory (CTIO); the MOSAIC-3 camera [46], formerly mounted on the Nicholas U. Mayall 4-meter Telescope at the Kitt Peak National Observatory; the 90Prime camera [47] on the Bok 2.3-meter telescope located on Kitt Peak; and the Wide-field Infrared Survey Explorer (WISE, [48]) satellite.
The three datasets described below are composed of these images and are processed with the pipeline Legacypipe. The DESI Legacy Imaging Survey DR9 described in Section 2.1 uses most of the images, excluding a few defective images. Images were also excluded because the selected region had reached the depth requirement. This set is used to mass produce Obiwan LRGs to study imaging systematics. The COSMOS Deep catalog in Section 2.3 utilizes the regions of the sky with many more exposures than a typical region and has much deeper photometry. This set is used to select the truth sample that Obiwan injects into the images. The COSMOS Repeats in Section 2.2 uses a large number of observations of the COSMOS field. The images are separated into ten distinct sets, each individually processed. These repeats are used to understand photometric scattering on DESI LRGs to define extended color cut for the injected Obiwan LRG-like sample.
@@ 2.1 DESI Legacy Imaging Survey DR9
We use and simulate data from the DR9 DESI Legacy Imaging Survey [^3] https://www.legacysurvey.org/dr9/description/ [33], which was used to obtain the sample that DESI uses for target selection and follow-up spectroscopy [49, 50, 51, 52, 53, 54, 55, 56]. We briefly describe these images here. This imaging survey consists of three optical bands: $g$, $r$, and $z$, and 4 infrared bands from WISE, $W1$, $W2$, $W3$, and $W4$. The optical band is separated into two parts, the South and the North, by a declination (Dec) of $32.375$ in the North Galactic Cap. The South consists of exposures taken from DECam [57], with a typical seeing of less than 1.3". The North part consists of images from the Beijing-Arizona Sky Survey [58] using the 90Prime camera for the $g$ and $r$ bands, and the Mosaic z-band Legacy Survey (MzLS) [59] for the $z$-band. The BASS survey has a typical seeing of less than 1.7", and MzLS has a less than 1.3" seeing. The pipeline transforms these images into a common grid with a pixel scale of 0.262 arcsec. These pixels are grouped into $3600\times 3600$ ‘bricks’, and image processing is performed at the individual brick level.
@@
The infrared bands are imaged by the WISE survey [60]. WISE is a space telescope whose orbits intersect with its Ecliptic poles ($|\textrm{Dec}|\approx 66$) while taking snapshots of the sky during its journey. Because of this orbiting plan, the WISE images have stable orientation in regions far from the Ecliptic poles. Image orientation can be very different during different visits for regions close to the Ecliptic poles. Because of this difference, it is much easier to measure regions far from the Ecliptic poles: The point spread function (PSF) of WISE galaxies is a stable image, and its profile is determined by the orientation of the WISE telescope [61]. Our work focuses only on the Southern part of the imaging survey because we do not have a truth input sample in the northern imaging area, and more details are explained in Section 2.3. Though WISE is deeper near the Ecliptic poles, these regions are more challenging to model due to the different image orientations in different visits; however, we avoid these problematic regions since our footprint does not cover that region. The WISE survey has four bandpasses, but only two bands ($W1$ and $W2$) are used in DESI target selection, so we will focus on only $W1$ and $W2$ in what follows. These two bands have a typical point-spread function width of 6.5". The WISE images we process are the unWISE coadded data products [62, 63], which use a 2.75" pixel scale. Figure 1 is a representative source with typical PSF size in all bands. It is located in a region observed by all three cameras in the optical bands. The first column shows the real images in optical and infrared bands. The second column shows galaxy models of the real images. The third column shows the galaxy stamp for the simulated galaxy. The fourth column is the “blob map”: sources within this map are treated as blended sources and will be fit simultaneously.
@@
Figure 1: Simulated targets in different bands and cameras, processed through the Legacypipe image reduction pipeline. The injected galaxy is located at the center of each image. The figures in the first column are the co-added images. The second column displays the galaxy/star profiles derived from fitting the image. The third column shows the simulated source injected into the images after convolution with the local PSF. The fourth column shows the contiguous area (known as a ‘blob’), where optical sources within the same blob are fit simultaneously (e.g., to resolve blending). The blobs are not used for WISE processing (see text for details) and, thus, are not shown for their images. The top rows show the WISE $W1$ and $W2$ band images, respectively. The third row shows the DECam image in the $g$- band, and the fourth row shows the $r$- band image from the BASS survey. The third row shows the $z$- band image from the MzLS survey.
@@ 2.2 COSMOS Repeats dataset
The COSMOS region has been imaged extensively by many instruments, including DECam. The COSMOS Repeats dataset we use is a set of catalogs covering $1\deg^{2}$ around (RA, Dec) = ($150.1\deg,2.2\deg$). This small region has a large number of exposures taken in all optical bands, which provides the ability to produce many independent subsets of data processed through Legacypipe and containing the same photometric sources. Ten such realizations were produced for the COSMOS repeats data. Gaussian noise was added to each image in each set so that the total galdepth [^4] It is defined as detection sensitivity of a round exponential galaxy with a radius of 0.45 arcsecond. in each set was equal to the DESI imaging requirements. The different image sets are grouped by seeing. The ten sets are labeled “set 0” through “set 9”, going from smallest to largest seeing (psfsize). The scatter in photometric measurements of sources common to each set of the COSMOS Repeats data thus provides a fully empirical determination of the expected variance in Legacypipe photometry in different seeing conditions. We simulate sources in the same sets of images with the same settings as COSMOS repeats catalog. We obtained ten sets of corresponding simulated Obiwan COSMOS repeat catalogs that mimic the scattering behavior of sources in real COSMOS repeat catalogs. Achieving consistency with the distributions recovered from the real data in the COSMOS Repeats is our main method of validating the LRG-Obiwan pipeline. The COSMOS Repeats data will also be used to determine our target selection for the extended color boundary that we draw our truth sample from and will be discussed in Section 4. It will also be used to study PSF error contributions in Section 7.3.